OpenStudy (anonymous):

A salmon is 2.0 meters away from a 0.55 meter tall waterfall. If the salmon jumps at a 32 degree angle, what is the minimum speed necessary to reach the fall? @Mathematics

OpenStudy (anonymous):

for minimum speed take the .55 meter as the max height put it in equation and solve, and also take 2.0 meters as the half horizontal range I think the formula for horizontal range is \(\Large \frac{u^2*\sin{2\theta}}{g}\) and for max height is \(\Large \frac{u^2*sin^2 \theta}{2g}\)

OpenStudy (anonymous):

Hmm, I've never seen those equations in my life haha. This is for a beginning physics course, where there's only 5 standard equations for motion, so do you happen to know how to solve it using those :)?

OpenStudy (anonymous):

break it down into the x and y componets for x component x-initial = 0 i use x to denote displacement x-final = 2.0m v-initial =v-final =? the x component of the velocity, thisis needed a = 0 t =? useful to find since it is independent of x or y direction for y component x-inital =0 x-final = 55.0m v-initial = ? this is the y component of velocity you need to find v-final =? not really needed to solve this question a = -9.8m/s^2 t=? use proper equations to solve for unknowns then use trig on v- initial components to solve for v-initial

OpenStudy (anonymous):

For both directions, you have 2 unknowns... So how do you solve it?